Qd = Q = 75 - 2P. +1 -2* 10 +0.1 * 100 = 66 - 2Pc -4.6 + 1Pc. (2.6) (2.7)

Using equation 2.6 and 2.7 from Dahl, how many tons of coal would buyers purchase if the price of coal were $50 per tons?

Pc is the price of coal. The answer is not 18.93 or 23.53
Answer #1

Qd = 66-2Pc

Qs = -4.6 + 1Pc

If Pc = $50 per ton , then Qd= 66-2(50)= -34 units

Qs = -4.6+ 1(50)= 45.4 units

Because if the price of coal were $50 per ton , then quantity demanded is negative , this implies that buyers would purchase 0 units of good at this price.